### v2

# Fully stochastic trust-region methods with Barzilai-Borwein steplengths

Abstract
We investigate stochastic gradient methods and stochastic counterparts of the Barzilai-Borwein steplengths and their application to finite-sum minimization problems. Our proposal is based on the Trust-Region-ish (TRish) framework introduced in [F. E. Curtis, K. Scheinberg, R. Shi, A stochastic trust region algorithm based on careful step normalization, Informs Journal on Optimization, 1, 2019]. The new framework, named TRishBB, aims to enhance the performance of TRish and at reducing the computational cost of the second-order TRish variant. We propose three different methods belonging to the TRishBB framework and present the convergence analysis for possibly nonconvex objective functions, considering biased and unbiased gradient approximations. Our analysis requires neither diminishing step-sizes nor full gradient evaluation. The numerical experiments in machine learning applications demonstrate the effectiveness of applying the Barzilai-Borwein steplength with stochastic gradients and show improved testing accuracy compared to the TRish method.
keywords:
finite-sum minimization, stochastic trust-region methods, Barzilai-Borwein method, machine learning.
MSC:
[2010]65K05 , 90C30
journal: Journal of Computational and Applied Mathematics
[…]
[label1]organization=Department of Industrial Engineering, University of Florence, addressline= Viale G.B Morgagni 40, Florence, 50134 (FI), country= Italy
1 Introduction
We consider the following optimization problem
$\min_{x\in\mathbb{R}^{n}}\,f(x)=\frac{1}{N}\sum_{i=1}^{N}\,f_{i}(x),$ (1)
which minimizes a finite-sum function $f$ with large $N$. We assume that each function $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$, $i\in\mathcal{N}\stackrel{{\scriptstyle\rm def}}{{=}}\{1,\ldots,N\}$, is continuously differentiable. Several important problems can be stated in this form, e.g., classification problems in Machine Learning (ML) or Deep Learning (DL), data fitting problems, and sample average approximations of an objective function given in the form of mathematical expectation. In the class of first-order optimization methods, Gradient Descent (GD) methods generate iterations of the form
$x_{k+1}=x_{k}-\mu_{k}\nabla f(x_{k}),$ (2)
for some positive scalar $\mu_{k}$ and $k\geq 0$. Convergence and efficiency of the GD method and more generally of gradient-based methods depend heavily on the sequence of the steplengths [1, 2, 3]. In [1] Barzilai and Borwein proposed to use a parameter $\mu_{k}$ which attempts to capture second-order information without computing the Hessian matrix of $f$ nor its inverse. In other words, the Barzilai-Borwein (BB) method can be regarded as a quasi-Newton method where the Hessian is approximated by $\mu_{k}^{-1}I$, $I\in\mathbb{R}^{n\times n}$ being the identity matrix, and $\mu_{k}$ minimizes the least-squares error of a secant equation involving the difference between two successive iterates and the difference between two successive gradients.
Computing the gradient $\nabla f(x_{k})$ to perform the update (2) is practically inconvenient in ML and DL applications where $f$ represents a loss function on a given dataset and $N$ is large. The Stochastic Gradient (SG) method [4, 5] and the mini-batch SG method [6] are iterative procedures where the full gradient $\nabla f(x_{k})$ is replaced by either a single gradient $\nabla f_{i}(x_{k})$, $i\in\mathcal{N}$ selected at random, or by a mini-batch of samples of the form $\nabla f_{\mathcal{N}_{k}}(x_{k})=\frac{1}{|\mathcal{N}_{k}|}\sum_{i\in\mathcal{N}_{k}}\nabla f_{i}(x_{k})$ where the set $\mathcal{N}_{k}\subseteq\mathcal{N}$ is selected at random and $|\mathcal{N}_{k}|$ denotes its cardinality, i.e., the mini-batch size. A limitation of such methods is the need to choose problem-dependent steplengths to guarantee convergence. An attempt to reduce such dependence was made in [7] where the steplengths depend on a Trust-Region-ish (TRish) rule and normalized steps are selected whenever the norm of the stochastic gradient is within a prefixed interval whose endpoints need to be tuned. The TRish method was designed for both first-order models [7, 8] and quadratic models [9] and, contrary to a standard trust-region approach, does not test for actual vs. predicted reduction and thus does not call for $f$-evaluations. Results from [7, 8] show that TRish can outperform a standard stochastic gradient approach and it was recently extended to equality-constrained problems [10].
The aim of this work is to combine the TRish methodology and stochastic BB steplengths in order to employ second-order information at a low computational cost; the selection of BB steplengths is ruled by TRish automatically. We devise the computation of BB steplengths in a stochastic setting and provide a theoretical analysis without requiring either diminishing steplengths nor unbiased stochastic gradients. In addition, we design three different procedures belonging to the above framework and show numerical results where our procedures enhance the performance of the first-order TRish method [9].
This paper is organized as follows. The remainder of this section is devoted to a review of the relevant literature providing background, motivation, and contribution to our work. We outline the proposed stochastic algorithms in Section 2 and present the theoretical analysis in Section 3. In Section 4 we present the implementation details, experimental setup, and an extensive evaluation of the method’s performance. Finally, in Section 5 we provide an overview of the obtained results and perspectives.
Notations
[…]
1.1 Related works and motivations
The use of BB steplengths combined with stochastic gradients calls for the definition of the stochastic counterparts of the standard BB steplengths and for convergence analysis in expectation or high probability of the resulting procedures. Several attempts have been made [11, 12, 13, 14, 15] and typically originate from the class of stochastic quasi-Newton methods [16, 17, 18, 19, 20, 21, 22].
In [11], BB steplengths were incorporated into the Stochastic Variance Reduction Gradient (SVRG) method [23] and into the SG method. The proposed methods are inner-outer loop-based frameworks. Since SVRG provides a full gradient at each outer loop, the BB steplength was computed at outer loops using the standard formula for the BB steplength; convergence of the resulting procedure was proved for strongly convex objective functions. On the other hand, single stochastic gradients $\nabla f_{i}$ were employed at each iteration in the SG method and, in the inner loop, a moving averaging rule over $N$ stochastic gradients at different points was computed; then, such average was used at each outer loop to evaluate a stochastic BB parameter. A theoretical support for this algorithm, denoted SGD-BB algorithm, was not provided. In [12], deterministic BB steplengths from [24], based on exact gradient and function evaluations, were used in the SVRG method; convergence was analyzed for strongly convex objective functions. In [13], stochastic BB steplengths were formed by a moving average rule on mini-batch gradients evaluated at different points and selected if they lied within an interval with endpoints tending to zero. Using diminishing stepsizes provides convergence due to well-known results for SG method but, in principle, BB steplengths may be discarded eventually since they are not supposed to tend to zero. A further method in which stochastic BB steplengths were formed and diminishing stepsizes were imposed is given in [14]; the distinguishing feature is that the mini-batch of sample is taken constant for a prefixed number of iterations with the aim to explore the underlying approximate Hessian spectrum. Finally, the paper [15] investigates on the use of BB steplengths combined with projected stochastic gradients for constrained optimization problems and with the objective function in the form of mathematical expectation.
The above mentioned papers show limitations such as use of diminishing stepsizes, use of increasing batch sizes $|{\cal{N}}_{k}|$, use of full gradient information, need for $f$-evaluations. Most of the theoretical results hold for convex or strongly convex objective functions. Our proposal overcomes the above drawbacks employing the TRish paradigm given in [7, 9] which allows to use a simple stochastic quadratic model inspired by the BB method and to select adaptively the step by a trust-region methodology not invoking $f$-evaluations. The algorithms proposed in this work are cheap to implement and supported by theoretical results for possibly nonconvex objective functions and for biased or unbiased gradients.
2 New Stochastic Trust-Region-ish Algorithms
Algorithm 1 TRishBB
1: Choose an initial iterate $x_{0}\in\mathbb{R}^{n}$ and a parameter $\mu_{0}>0$.
2: Choose $(\alpha,\gamma_{1},\gamma_{2})$ such that $\alpha>0$, $0<\gamma_{2}\leq\gamma_{1}<\infty$ and $\mu_{min}>0$.
3:for $k=0,1,\ldots$ do
4:  Set $H_{k}=\mu_{k}^{-1}I$.
5:  Compute a stochastic gradient $g_{k}$.
6:  Compute the solution $p_{k}$ of the trust-region subproblem
$\min_{p\in\mathbb{R}^{n}}m_{k}(p)=g_{k}^{T}p+\frac{1}{2}p^{T}H_{k}p\quad\text{s.t.}\quad\|p\|\leq\Delta_{k},$ (3)
where
$\Delta_{k}=\begin{cases}\alpha\gamma_{1}\|g_{k}\|&\quad\text{if }\|g_{k}\|\in\left[0,\frac{1}{\gamma_{1}}\right)\\ \alpha&\quad\text{if }\|g_{k}\|\in\left[\frac{1}{\gamma_{1}},\frac{1}{\gamma_{2}}\right]\\ \alpha\gamma_{2}\|g_{k}\|&\quad\text{if }\|g_{k}\|\in\left(\frac{1}{\gamma_{2}},\infty\right)\end{cases}.$ (4)
7:  Set $x_{k+1}=x_{k}+p_{k}$.
8:  Form a parameter $\mu_{k+1}\geq\mu_{min}$.
9:end for
In this section we present our trust-region-ish algorithm employing stochastic BB steplengths. First, we provide a brief overview of the BB gradient method in the deterministic setting [1, 25, 26, 27], second we describe the main features of the method denoted as TRishBB, finally we present three algorithms belonging to the TRishBB frameworks.
At $k$-th iteration, the standard BB method minimizes the second-order model $m_{k}(p)=\nabla f(x_{k})^{T}p+\frac{1}{2}p^{T}H_{k}p$ where $H_{k}=\mu_{k}^{-1}I$ for some positive $\mu_{k}$; thus the iterates have the form (2). Regarding the computation of the steplength, given $x_{k+1}$, the main idea of the BB method is to form the parameter $\mu_{k+1}^{\text{BB}}$ which solves
$\mu_{k+1}^{\text{BB}}=\text{argmin}_{\mu}\ \|\mu^{-1}s_{k}-y_{k}\|^{2}=\frac{s_{k}^{T}s_{k}}{s_{k}^{T}y_{k}},$ (5)
with $s_{k}=x_{k+1}-x_{k}$, and $y_{k}=\nabla f(x_{k+1})-\nabla f(x_{k})$ commonly denoted as correction pair. If $f$ is strictly convex, then the scalar $\mu_{k+1}^{\text{BB}}$ is positive.
If $f$ is twice continuously differentiable then $\mu_{k+1}^{\text{BB}}$ is related to the spectrum of the Hessian matrix of $f$ as follows. Given a symmetric matrix $A\in\mathbb{R}^{n\times n}$, let $q(A,v)$ be the Rayleigh quotient associated to the vector $v\in\mathbb{R}^{n}$; in addition, let $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote the smallest and largest eigenvalue of $A$, respectively. Then, by the mean value theorem [28, Th. 3.1.3], we have
[…]
In the BB method, $H_{k+1}=\mu_{k+1}^{-1}I$ is formed using $|\mu_{k+1}^{\text{BB}}|$ and thresholding $|\mu_{k+1}^{\text{BB}}|$ by means of two predefined scalars $0<\mu_{min}<\mu_{max}$, i.e., setting $\mu_{k+1}=\max\{\mu_{min},\min\{|\mu_{k+1}^{\text{BB}}|,\mu_{max}\}$ to prevent excessively small or large values; in the case of convex problems it trivially holds $\mu_{k+1}=\max\{\mu_{min},\min\{\mu_{k+1}^{\text{BB}},\mu_{max}\}$. We refer to [27] for a survey on both practical rules for fixing $\mu_{k+1}$ and step acceptance rules in the deterministic convex and nonconvex settings.
Now we analyze Algorithm 1 which is based on the second-order version of TRish in [9] and specialized to the case where the Hessian matrix is approximated by the diagonal matrix $H_{k}=\mu_{k}^{-1}I$, $\forall k$. Our goal is to use steplengths $\mu_{k}$’s inspired by the BB method and to enhance the first-order version of TRish [7] where $H_{k}=0$ for all $k\geq 0$. However, since stochastic gradients are available, the correction pair $(s_{k},y_{k})$ is not uniquely defined and $\mu_{k+1}^{\text{BB}}$ in (5) is a random parameter. The choice of the pair $(s_{k},y_{k})$ is crucial for the practical behaviour of the method and we postpone the discussion of three options for the definition of $(s_{k},y_{k})$ and $\mu_{k+1}$ to Subsections 2.1-2.3.
At the generic $k$-th iteration of TRishBB, the stochastic gradient $g_{k}$ is computed. Then, given the trust-region radius $\Delta_{k}$ and the scalar $\mu_{k}>\mu_{min}$, we set $H_{k}=\mu_{k}^{-1}I$ and let (3) be the trust-region problem with strictly convex quadratic model. The trust-region solution is cheap to compute and takes the form
$p_{k}=\begin{cases}-\mu_{k}g_{k}&\quad\text{if }\|\mu_{k}g_{k}\|<\Delta_{k},\\ \displaystyle-\frac{\Delta_{k}}{\|g_{k}\|}g_{k}&\quad\text{otherwise. }\end{cases}$ (7)
In fact, when $p_{k}$ takes the form $p_{k}=-H_{k}^{-1}g_{k}=-\mu_{k}g_{k}$, it holds $\|p_{k}\|<\Delta_{k}$ and $p_{k}$ is the unconstrained minimizer of the model $m_{k}$. We refer to such a step as the unconstrained solution of (3). Otherwise, the solution lies on the boundary of the trust-region and it is the constrained Cauchy point; we will refer to this step as the constrained solution. In both cases, $p_{k}$ is a step along $(-g_{k})$ and the cost for solving the trust-region problem is low. On the contrary, using a second-order trust-region methods and a general $H_{k}$ requires a linear algebra phase such as the application of the conjugate gradient method [28, §7.5.1].
The updating rule (4) for $\Delta_{k}$ is a distinctive feature of TRish. It is ruled by three positive and predefined scalars, $\alpha$, $\gamma_{1}$ and $\gamma_{2}$ and the magnitude of the norm of $g_{k}$. The insight of this rule is given in [7, §2]; briefly, adopting the standard rule, i.e., setting $\Delta_{k}=\alpha$, and thus normalizing $(-g_{k})$ if the trust-region solution is constrained, may cause the algorithm to compute ascent directions more likely than descent directions and then fail to progress in expectation. We note that in [7, 9] the algorithm admits the use of positive sequences $\{\alpha_{k}\}$, $\{\gamma_{1,k}\}$ and $\{\gamma_{2,k}\}$ but here we focus on the choice of constant parameters.
Due to (4) and (7), the constrained trust-region solution has the form
$p_{k}=\begin{cases}-\alpha\gamma_{1}g_{k}&\quad\text{if }\|g_{k}\|\in\left[0,\frac{1}{\gamma_{1}}\right)\\ -\displaystyle\alpha\frac{g_{k}}{\|g_{k}\|}&\quad\text{if }\|g_{k}\|\in\left[\frac{1}{\gamma_{1}},\frac{1}{\gamma_{2}}\right]\\ -\alpha\gamma_{2}g_{k}&\quad\text{if }\|g_{k}\|\in\left(\frac{1}{\gamma_{2}},\infty\right)\end{cases},$ (8)
analogously to the first-order TRish method in [7], whereas the unconstrained trust-region solution is $p_{k}=-\mu_{k}g_{k}$ when
$\mu_{k}<\begin{cases}\alpha\gamma_{1}&\quad\text{if }\|g_{k}\|\in\left[0,\frac{1}{\gamma_{1}}\right)\\ \displaystyle\frac{\alpha}{\|g_{k}\|}&\quad\text{if }\|g_{k}\|\in\left[\frac{1}{\gamma_{1}},\frac{1}{\gamma_{2}}\right]\\ \alpha\gamma_{2}&\quad\text{if }\|g_{k}\|\in\left(\frac{1}{\gamma_{2}},\infty\right)\end{cases}.$ (9)
Importantly, once the step $p_{k}$ is formed, the new iterate $x_{k+1}$ is computed without performing the standard test of trust-region methods on predicted vs. actual reduction [28]; thus the Algorithm TRishBB does not require the evaluation of the objective function $f$.
We conclude this section noting that each step $p_{k}$ taken by TRishBB is along $(-g_{k})$ as in a first-order method but the unconstrained trust-region solution aims to exploit some second-order information via $\mu_{k}$.
In the following subsections we explore three ways to form stochastic BB steplengths; each of them provides a variant of the general Algorithm TRishBB.
2.1 TRishBB_v1
Algorithm 2 TRishBB_v1 (Iteration $k$)
1: Given $x_{k}\in\mathbb{R}^{n}$, $(\alpha,\gamma_{1},\gamma_{2})$, $\mu_{k}>0$, $0<\mu_{min}<\mu_{max}$, $m\geq 1$.
2: Choose $\mathcal{N}_{k}\subseteq\mathcal{N}$, compute $g_{k}=\nabla f_{\mathcal{N}_{k}}(x_{k})$ and set $H_{k}=\mu_{k}^{-1}I$.
3: Obtain $p_{k}$ from (7) and set $x_{k+1}=x_{k}+p_{k}$.
4:if $\text{{mod}}(k,m)=0$ then
5:  $s_{k}=p_{k}$ and $y_{k}=\nabla f_{\mathcal{N}_{k}}(x_{k+1})-g_{k}$
6:  $\mu_{k+1}=\max\left\{\mu_{min},\min\left\{\left|\frac{s_{k}^{T}s_{k}}{s_{k}^{T}y_{k}}\right|,\mu_{max}\right\}\right\}$
7:else
8:  $\mu_{k+1}=\mu_{k}$
9:end if
The displayed Algorithm TRishBB_v1 is a straightforward implementation of the BB method employing stochastic gradients and guidelines from stochastic quasi-Newton methods. Specifically, once the trust-region solution $p_{k}$ in (7) is computed, the successive steplength $\mu_{k+1}$ is formed using the correction pair $(s_{k},y_{k})$ at line 5 of this algorithm. The vector $s_{k}$ coincides with $p_{k}$, i.e., it is the step between $x_{k+1}$ and $x_{k}$ whereas the vector $y_{k}$ is the difference between the mini-batch gradients $\nabla f_{\mathcal{N}_{k}}(x_{k+1})$ and $\nabla f_{\mathcal{N}_{k}}(x_{k})$. We underline that using the same set of indexes $\mathcal{N}_{k}$ in the computation of $y_{k}$ is suggested in [29, 16, 17] and implies that (6) holds with the matrix $\int_{0}^{1}\nabla^{2}f_{\mathcal{N}_{k}}(x_{k}+ts_{k})$. Thus, the BB steplength inherits spectral information as long as it is positive and $f$ is twice continuously differentiable. Following a heuristic common rule, see e.g., [30], given an integer $m\geq 1$, the parameter $\mu_{k+1}$ is updated every $m$ iterations.
Compared to TRish, the extra cost of this algorithm amounts to the computation of $\nabla f_{\mathcal{N}_{k}}(x_{k+1})$ and two scalar products to evaluate $\mu_{k+1}$.
Algorithm 3 TRishBB_v2 (Iteration $k$)
1: Given $x_{k}\in\mathbb{R}^{n}$, $(\alpha,\gamma_{1},\gamma_{2})$, $\mu_{k}>0$, $\bar{\mu}>0$, $0<\mu_{min}<\mu_{max}$, $m\geq 1$, $\beta=\frac{m-1}{m}$, $\eta\in(0,1)$, $\bar{x}_{old}$, $\bar{g}_{old}$, $\bar{g}$  (if $k=0$, then $\bar{x}_{old}=x_{0}$, $\bar{g}_{old}=0$, $\bar{g}=0$, and $\bar{\mu}={\mu}_{0}$):
2: Choose $\mathcal{N}_{k}\subseteq\mathcal{N}$, compute $g_{k}=\nabla f_{\mathcal{N}_{k}}(x_{k})$ and set $H_{k}=\mu_{k}^{-1}I$;
3: Obtain $p_{k}$ from (7) and set $x_{k+1}=x_{k}+p_{k}$;
4: Compute $\bar{g}=\beta\bar{g}+(1-\beta)g_{k}$;
5:if $k>0$  and  $\text{{mod}}(k,m)=0$ then
6:  $\bar{x}=x_{k+1}$
7:  $s_{k}=\bar{x}-\bar{x}_{old}$ and $y_{k}=\bar{g}-\bar{g}_{old}$
8:   $\widehat{\mu}_{k}=\frac{1}{m}\left|\frac{s_{k}^{T}s_{k}}{s_{k}^{T}y_{k}}\right|$, and $\bar{\mu}=\eta\bar{\mu}+(1-\eta)\widehat{\mu}_{k}$
9:   $\mu_{k+1}=\max\{\mu_{min},\min\{\bar{\mu},\mu_{max}\}\}$
10:  $\bar{x}_{old}=\bar{x}$, $\bar{g}_{old}=\bar{g}$, $\bar{g}=0$
11:else
12:  $\mu_{k+1}=\mu_{k}$
13:end if
2.2 TRishBB_v2
In Algorithm TRishBB_v1 the pair $(s_{k},y_{k})$ contains information from one single iteration and $y_{k}$ is obtained using stochastic gradients, possibly evaluated on a small sample ${\mathcal{N}_{k}}$. On the contrary, Algorithm TRishBB_v2 forms a correction pair that accumulates information over a prefixed number $m$ of iterations and is described below.
The vector $s_{k}$ is the difference between two, not consecutive, iterates; the indexes of such iterates differ by $m$, see lines 7 and 10 of the algorithm. The vector $y_{k}$ is the difference between vectors computed via a moving average rule which accumulates stochastic gradients over $m$ iterations, see lines 4, 7 and 10 of the algorithm; in particular $\bar{g}$ is formed over the last $m$ iterations while $\bar{g}_{old}$ is the analogous vector formed in the previous cycle of $m$ iterations. The parameter $\beta=(m-1)/m$ is fixed in accordance with [31] to avoid hyper-parameter tuning.
The form of $s_{k}$ and $y_{k}$ and $\widehat{\mu}_{k}$ is borrowed from the stochastic gradient descent algorithm with BB step size (SGD-BB) algorithm proposed in [11] where $|\mathcal{N}_{k}|=1$, $\forall k$, and $m=N$, i.e., $N$ single gradient evaluations are performed before updating the steplength. In TRishBB_v2 we allow $|\mathcal{N}_{k}|>1$; letting $N_{b}=\lfloor N/|\mathcal{N}_{k}|\rfloor$ be (roughly) the number of disjoint mini-batches of size $|\mathcal{N}_{k}|$ in $\mathcal{N}$ and $m=N_{b}$, the cost for computing $\bar{g}$ through $m$ iterations is approximately the cost for one full gradient evaluation.
The steplength $\mu_{k+1}$ is updated every $m$ iterations and computed using accumulated information, see lines 8 and 9 of the algorithm. Given $\eta\in(0,1)$, the scalar $\bar{\mu}$ is computed as the convex combination of the current stochastic BB steplength $\widehat{\mu}_{k}$ and the current accumulated parameter $\bar{\mu}$. The larger $\eta$ is, the higher the penalization to $\widehat{\mu}_{k}$ is and this may soften the effect of intrinsic noise arising from stochastic gradients.
Compared to TRish, the extra cost of this algorithm amounts to two scalar products to evaluate $\mu_{k+1}$; in terms of storage, three vectors of size $n$ are required.
Algorithm 4 TRishBB_v3 (Iteration $k$)
1: Given $x_{k}\in\mathbb{R}^{n}$, $(\alpha,\gamma_{1},\gamma_{2})$, $\mu_{k}>0$, $\bar{\mu}>0$, $0<\mu_{min}<\mu_{max}$, $m\geq 1$, $m_{F}\in\mathbb{N}$ $F_{k}\in\mathbb{R}^{n\times q}$ with $q\leq m_{F}$, $x_{avg}$, and $\bar{x}_{old}$,   (if $k=0$, then $F_{0}=[\,]$, $x_{avg}=0$, and $\bar{x}_{old}=0$, and $\bar{\mu}={\mu}_{0}$).
[…]
3: Choose $\mathcal{N}_{k}\subseteq\mathcal{N}$, compute $g_{k}=\nabla f_{\mathcal{N}_{k}}(x_{k})$ and set $H_{k}=\mu_{k}^{-1}I$.
4: Obtain $p_{k}$ from (7) and set $x_{k+1}=x_{k}+p_{k}$.
5:if $k\leq m_{F}$ then
6:  Store $g_{k}$ in $k$-th column of $F_{k}$.
7:else
8:  Remove the first column of $F_{k}$, and store $g_{k}$ in its $m_{F}$-th column.
9:end if
10:if $\text{{mod}}(k,m)=0$ then
11:  $\bar{x}=\frac{x_{avg}}{m}$, and $x_{avg}=0$
12:  if $k>0$ then
13:   ${s}_{k}=\bar{x}-\bar{x}_{old}$, and ${y}_{k}=\frac{1}{|F_{k}|}F_{k}(F_{k}^{T}s_{k})$
14:    $\widehat{\mu}_{k}=\frac{1}{m}\left|\frac{s_{k}^{T}s_{k}}{s_{k}^{T}y_{k}}\right|$, and $\bar{\mu}=\eta\bar{\mu}+(1-\eta)\widehat{\mu}_{k}$;
15:    $\mu_{k+1}=\max\{\mu_{min},\min\{\bar{\mu},\mu_{max}\}\}$;
16:  end if
17:  $\bar{x}_{old}=\bar{x}$
18:else
19:  $\mu_{k+1}=\mu_{k}$
20:end if
2.3 TRishBB_v3
Algorithm TRishBB_v3 is originated by the arguments in [32] to avoid both the potentially harmful effects of the gradient difference when $\|s_{k}\|$ is small and the noise in stochastic gradient estimates. In [32], given $s_{k}=\bar{x}-\bar{x}_{old}$ for some $\bar{x},\,\bar{x}_{old}$, the vector $y_{k}$ is defined as $y_{k}=\nabla^{2}f_{\widetilde{\mathcal{N}_{k}}}(\bar{x})s_{k}$ where $\nabla^{2}f_{\widetilde{\mathcal{N}_{k}}}(\bar{x})$ is a subsampled Hessian approximation and $|\widetilde{\mathcal{N}_{k}}|$ is supposed to be larger than the batch size used for gradients. If the objective function $f$ is the cross-entropy loss function or the least-squares function, then $\nabla^{2}f_{\widetilde{\mathcal{N}_{k}}}(\bar{x})$ can be approximated by the empirical Fisher Information Matrix (eFIM) [33, §11] defined as
$\frac{1}{|\widetilde{\mathcal{N}_{k}}|}\sum_{i\in\widetilde{\mathcal{N}_{k}}}\nabla f_{i}(\bar{x})\nabla f_{i}(\bar{x})^{T}.$
In order to reduce the cost, the eFIM was further simplified by introducing the accumulated Fisher Information Matrix (aFIM) [34, 35]. In Algorithm TRishBB_v3, we update $s_{k}$ and $y_{k}$ every $m$ iterations, and $y_{k}$ is computed using the aFIM and the first-in, first-out (FIFO) approach [34, 35]. In particular, as displayed in line 13 of the algorithm, $s_{k}$ is formed using the averages $\bar{x}$ and $\bar{x}_{old}$; $\bar{x}$ is the average of the last $m$ iterates, and $\bar{x}_{old}$ is the average of the previous $m$ iterates, see lines 2 and 11 of the algorithm. Average iterates are meant to represent more stable approximations than a single iterate for computing the difference $s_{k}$. Concerning $y_{k}$, let $F_{k}$ be the limited memory matrix whose columns contain at most the last $m_{F}$ computed stochastic gradients for some positive integer $m_{F}$. Then, the curvature vector $y_{k}$ is defined as ${y}_{k}=\frac{1}{|F_{k}|}F_{k}(F_{k}^{T}s_{k})$, see line 13 of the algorithm, where $|F_{k}|$ denotes the number of columns of $F_{k}$. Since $F_{k}F_{k}^{T}$ is the sum of $|F_{k}|$ rank-one matrices, (6) implies
$\widehat{\mu}_{k}=\frac{|F_{k}|}{m}\frac{1}{\sum_{t=k-|F_{k}|+1}^{k}q(\nabla f_{\mathcal{N}_{t}}(x_{t})(\nabla f_{\mathcal{N}_{t}}(x_{t}))^{T},{s}_{k})},$
where $\nabla f_{\mathcal{N}_{t}}(x_{t})(\nabla f_{\mathcal{N}_{t}}(x_{t}))^{T}$ is a rank-one matrix. Finally, the update of $\mu_{k}$ is performed every $m$ iterations and it is the same as in TRishBB_v2.
Compared to TRish, the extra cost of this algorithm amounts to $m_{F}+2$ scalar products to evaluate $\mu_{k+1}$; in terms of storage, two vectors of size $n$ and one matrix of size $n\times m_{F}$ are required.
3 Convergence Analysis
We present the convergence analysis of TRishBB considering the use of both biased and unbiased stochastic gradients. Further, we investigate the case where the function $f$ satisfies the Polyak-Lojasiewicz (PL) condition and the case where $f$ is a general nonconvex function.
The assumption of the objective function made in our analysis and the PL condition are presented below.
Assumption 3.1.
Function $f:\mathbb{R}^{n}\to\mathbb{R}$ is continuously differentiable, bounded below by $f_{*}$, and there exists a positive $L$ such that
$f(x)\leq f(y)+\nabla f(y)^{T}(x-y)+\frac{L}{2}\|x-y\|^{2}\ \ \text{for all}\ \ x,y\in\mathbb{R}^{n}.$ (10)
Assumption 3.2.
For any $x\in\mathbb{R}^{n}$, the Polyak-Lojasiewicz (PL) condition holds with $c\in(0,\infty)$, i.e.,
$2c(f(x)-f_{*})\leq\|\nabla f(x)\|^{2}\ \ \text{for all}\ \ x\in\mathbb{R}^{n}.$ (11)
In our analysis we characterize the iterations of TRishBB by three cases: “Case $:\|g_{k}\|\in\left[0,\tfrac{1}{\gamma_{1}}\right)$, “Case 2”: $\|g_{k}\|\in\left[\tfrac{1}{\gamma_{1}},\tfrac{1}{\gamma_{2}}\right]$, and “Case 3”: $\|g_{k}\|\in\left(\tfrac{1}{\gamma_{2}},\infty\right)$. We denote by $C_{i,k}$ the events where Case $i$ occurs, for $i\in\{1,2,3\}$.
3.1 Analysis with Biased Gradients
The results presented in this subsection concern biased gradients satisfying the following assumption.
Assumption 3.3.
There exist positive $\omega,\,M_{1},M_{2}$ such that the stochastic gradient $g_{k}$ satisfy
$\nabla f(x_{k})^{T}\mathbb{E}_{k}[g_{k}]\geq\omega\|\nabla f(x_{k})\|^{2},$ (12)
$\mathbb{E}_{k}[\|g_{k}\|^{2}]\leq M_{1}+M_{2}\|\nabla f(x_{k})\|^{2}.$ (13)
The inequality (13) implies that
$\mathbb{P}_{k}[{\cal{E}}_{k}]\mathbb{E}_{k}[\nabla f(x_{k})^{T}g_{k}|{\cal{E}}_{k}]\leq h_{1}+h_{2}\|\nabla f(x_{k})\|^{2},$ (14)
with $h_{1}=\frac{1}{2}\sqrt{M_{1}}$ and $h_{2}=h_{1}+\sqrt{M_{2}}$; see [7, Lemma 2].
We also make the following assumption of the parameter $\mu_{min}$.
Assumption 3.4.
Given positive $\mu_{min}$, for all $k\geq 0$, it holds
$\mu_{min}\leq\gamma_{2}\alpha.$
The results presented are based on the analysis developed by [7]. The following two lemma are technical results. As for the notation, for the events of taking the unconstrained and the constrained solutions (7) we use the notations $S_{BB}$ and $\overline{S_{BB}}$, respectively. Moreover, for the events $C_{i,k}$, $i\in\{1,2,3\}$, that occur within each defined case we set ${C_{i,k}^{BB}}\stackrel{{\scriptstyle\rm def}}{{=}}C_{i,k}\cap S_{BB}$ and $\overline{C_{i,k}^{BB}}\stackrel{{\scriptstyle\rm def}}{{=}}C_{i,k}\cap\overline{S_{BB}}$. Finally, we denote with ${\cal{E}}_{k}$ the event that $\nabla f(x_{k})^{T}g_{k}\geq 0$ and with $\overline{{\cal{E}}_{k}}$ the event that $\nabla f(x_{k})^{T}g_{k}<0$.
Lemma 3.1.
Under Assumptions 3.1 and 3.4, the iterates generated by Algorithm TRishBB satisfy
[…]
for all $k$, where ${\cal{E}}_{k}$ is the event that $\nabla f(x_{k})^{T}g_{k}\geq 0$.
Proof.
See Section B.1. ∎
Lemma 3.2.
Under Assumptions 3.1,3.3, and 3.4, the iterates generated by Algorithm TRishBB satisfy
$\mathbb{E}_{k}[f(x_{k+1})]-f(x_{k})\leq(\gamma_{1}\alpha-\mu_{min})h_{1}+\frac{L}{2}\gamma_{1}^{2}\alpha^{2}M_{1}$ (16)
$-\gamma_{1}\alpha\left(\omega-\frac{L}{2}\gamma_{1}M_{2}\alpha\right)\|\nabla f(x_{k})\|^{2}+(\gamma_{1}\alpha-\mu_{min})h_{2}\|\nabla f(x_{k})\|^{2},$
for all $k$ with $h_{1},h_{2}$ given in (14).
Proof.
It is a direct consequence of Assumption 3.3, Lemma 3.1 and (14). ∎
For the rest of analysis, we define
$\theta_{1}=\frac{1}{2}\omega\gamma_{1}-{h_{2}}(\gamma_{1}-\rho\gamma_{2}),\quad\theta_{2}=h_{1}(\gamma_{1}\alpha\ -\mu_{min})+\frac{1}{2}\gamma_{1}^{2}LM_{1}\alpha^{2},$ (17)
where $\rho=\frac{\mu_{min}}{\alpha\gamma_{2}}\in(0,1]$, $h_{1}$ and $h_{2}$ are given in (14), $\omega$ and $M_{1}$ are given in Assumption 3.3. Note that $\theta_{1}$ is positive for any choice of $(\gamma_{1},\gamma_{2})$ in the case $h_{2}-\frac{1}{2}\omega\leq 0$, while condition
$\frac{\gamma_{1}}{\gamma_{2}}<\frac{\rho h_{2}}{h_{2}-\frac{1}{2}\omega}$ (18)
ensures $\theta_{1}>0$ in the case $h_{2}-\frac{1}{2}\omega>0$. Further, note that $\theta_{2}$ is positive by Assumption 3.4.
First we consider the case where the objective function satisfy the Polyak-Lojasiewicz condition (3.2), then we analyze the general case.
Theorem 3.1.
Under Assumptions 3.13.4, suppose that (18) holds in case $h_{2}-\frac{1}{2}\omega>0$ and that
$\alpha\leq\min\left\{\frac{\omega}{\gamma_{1}LM_{2}},\frac{1}{2c\theta_{1}}\right\},$ (19)
with $h_{2}$ given in (14) and $\theta_{1}$ given in (17). Then, the expected optimality gap of the iterates generated by Algorithm TRishBB satisfies
$\mathbb{E}[f(x_{k+1})]-f_{*}\xrightarrow{k\rightarrow\infty}\frac{\theta_{2}}{2c\alpha\theta_{1}},$ (20)
where $\theta_{2}$ is given in (17).
Proof.
By (19) and (16), it follows that
$\mathbb{E}_{k}[f(x_{k+1})]-f(x_{k})\leq$ $\ -\alpha\theta_{1}\|\nabla f(x_{k})\|^{2}+\theta_{2},$
for all $k$ and $\theta_{1},\,\theta_{2}$ defined in (17). Therefore, (11) implies that
$\mathbb{E}_{k}[f(x_{k+1})]-f(x_{k})$ $\leq-2c\alpha\theta_{1}(f(x_{k})-f_{*})+\theta_{2}.$ (21)
Adding and subtracting $f_{*}$ on the left-hand side, taking total expectations, and rearranging give
$\mathbb{E}[f(x_{k+1})]-f_{*}$ $\leq(1-2c\alpha\theta_{1})(\mathbb{E}[f(x_{k})]-f_{*})+\theta_{2}$ (22)
$=\frac{\theta_{2}}{2c\alpha\theta_{1}}+(1-2c\alpha\theta_{1})(\mathbb{E}[f(x_{k})]-f_{*})+\theta_{2}-\frac{\theta_{2}}{2c\alpha\theta_{1}}$
$=\frac{\theta_{2}}{2c\alpha\theta_{1}}+(1-2c\alpha\theta_{1})\left(\mathbb{E}[f(x_{k})]-f_{*}-\frac{\theta_{2}}{2c\alpha\theta_{1}}\right).$
Since $1-2c\alpha\theta_{1}\in(0,1)$, this represents a contraction inequality. Applying the result repeatedly through iteration $k$, one obtains the desired result. ∎
Theorem 3.2.
Under Assumptions 3.1, 3.3 and 3.4, suppose that (18) holds in case $h_{2}-\frac{1}{2}\omega>0$ and that
$\alpha\leq\frac{\omega}{\gamma_{1}LM_{2}},$ (23)
with $h_{2}$ given in (14) and $\theta_{1}$ given in (17). Then the iterates generated by Algorithm TRishBB satisfy
$\sum_{k=0}^{K}\mathbb{E}\left[\frac{1}{K}\|\nabla f(x_{k})\|^{2}\right]\xrightarrow{K\rightarrow\infty}\frac{\theta_{2}}{\alpha\theta_{1}},$ (24)
where $\theta_{2}$ is given in (17).
Proof.
By (23) and (16), it follows that
$\mathbb{E}_{k}[f(x_{k+1})]-f(x_{k})\leq$ $\ -\alpha\theta_{1}\|\nabla f(x_{k})\|^{2}+\theta_{2},$
for all $k$ and $\theta_{1},\,\theta_{2}$ defined in (17). Therefore, taking total expectation yields
$\mathbb{E}[\|\nabla f(x_{k})\|^{2}]$ $\leq\frac{1}{\alpha\theta_{1}}\left(\mathbb{E}[f(x_{k})]-\mathbb{E}[f(x_{k+1})]\right)+\frac{\theta_{2}}{\alpha\theta_{1}}.$ (25)
Summing both sides for $k\in\{1,2,\ldots,K\}$ gives
$\sum_{k=0}^{K}\mathbb{E}[\|\nabla f(x_{k})\|^{2}]$ $\leq\frac{1}{\alpha\theta_{1}}\left(f(x_{0})-\mathbb{E}[f(x_{K+1})]\right)+\frac{K\theta_{2}}{\alpha\theta_{1}}$
$\leq\frac{f(x_{0})-f_{*}}{\alpha\theta_{1}}+\frac{K\theta_{2}}{\alpha\theta_{1}}.$
Dividing by $K$ concludes the proof. ∎
3.2 Analysis with Unbiased Gradients
In this subsection we assume that $g_{k}$ is unbiased estimator of the true gradient with bounded variance. Such assumption is summarized below.
Assumption 3.5.
For all $k$, $g_{k}$ is an unbiased estimator of the gradient $\nabla f(x_{k})$, i.e., $\mathbb{E}_{k}[g_{k}]=\nabla f(x_{k})$. For all $k$, there exist a positive constant $M_{g}$ such that
$\mathbb{E}_{k}[\|\nabla f(x_{k})-g_{k}\|^{2}]\leq M_{g}.$ (26)
Since $g_{k}$ is unbiased, it follows.
$\mathbb{E}_{k}[\|\nabla f(x_{k})-g_{k}\|^{2}]$ $=\mathbb{E}_{k}[\|\nabla f(x_{k})\|^{2}-2\nabla f(x_{k})^{T}g_{k}+\|g_{k}\|^{2}]$ (27)
$=-\|\nabla f(x_{k})\|^{2}+\mathbb{E}_{k}[\|g_{k}\|^{2}].$
Hence Assumption (3.3) is satisfied with $\omega=1$, $M_{1}=M_{g}$, $M_{2}=1$ and the convergence results of Section 3.1 hold. Now we further characterize the convergence without imposing bounds, such as inequality (18), on $\gamma_{1}$, $\gamma_{2}$. The first lemma is a technical result.
Lemma 3.3.
Suppose that Assumption 3.1 holds and
$0<\alpha\leq\frac{\gamma_{2}}{8\gamma_{1}^{2}L},\qquad\mu_{min}\geq\frac{4\gamma_{1}\alpha}{5}.$ (28)
Then, the $k$-th iterate generated by Algorithm TRishBB satisfies
$\mathbb{E}_{k}[f(x_{k+1})]-f(x_{k})\leq-\frac{1}{16}\gamma_{2}\alpha\mathbb{E}_{k}[\|g_{k}\|^{2}]+\frac{\gamma_{1}^{2}}{\gamma_{2}}\alpha\mathbb{E}_{k}[\|\nabla f(x_{k})-g_{k}\|^{2}],\qquad$ (29)
$\mathbb{E}_{k}[f(x_{k+1})]-f(x_{k})\leq\ -\frac{1}{16}\gamma_{2}\alpha\|\nabla f(x_{k})\|^{2}+\alpha\left(\frac{\gamma_{1}^{2}}{\gamma_{2}}-\frac{1}{16}\gamma_{2}\right)M_{g}.\qquad$ (30)
Proof.
See Section B.2. ∎
In case the objective function satisfies the Polyak-Lojasiewicz condition (3.2), the expected optimality gap is bounded above by a sequence that converges linearly to a constant proportional to $M_{g}/c$.
Theorem 3.3.
Suppose that Assumptions 3.1, 3.2 and 3.5 hold and that
$0<\alpha<\min\left\{\frac{\gamma_{2}}{8\gamma_{1}^{2}L},\frac{8}{\gamma_{2}c}\right\},\qquad\mu_{min}\geq\frac{4\gamma_{1}\alpha}{5}.$ (31)
Then, the sequence generated by Algorithm TRishBB satisfies
$\mathbb{E}[f(x_{k+1})]-f_{*}\xrightarrow{k\rightarrow\infty}\frac{8\theta_{3}M_{g}}{\gamma_{2}c},$ (32)
for all $k$, where $\theta_{3}=\left(\frac{\gamma_{1}^{2}}{\gamma_{2}}-\frac{\gamma_{2}}{16}\right)$.
Proof.
By (30) and (11) we obtain
$\mathbb{E}_{k}[f(x_{k+1})]-f(x_{k})$ $\leq-\frac{1}{8}\gamma_{2}\alpha c(f(x_{k})-f_{*})+\alpha\theta_{3}M_{g}.$ (33)
Taking total expectation, for all $k$ we have
$\mathbb{E}[f(x_{k+1})]-f_{*}$ $\leq\frac{8\theta_{3}M_{g}}{\gamma_{2}c}+\left(1-\frac{1}{8}\gamma_{2}\alpha c\right)(\mathbb{E}[f(x_{k})]-f_{*})+\alpha\theta_{3}M_{g}-\frac{8\theta_{3}M_{g}}{\gamma_{2}c}$ (34)
$=\frac{8\theta_{3}M_{g}}{\gamma_{2}c}+\left(1-\frac{1}{8}\gamma_{2}\alpha c\right)\left(\mathbb{E}[f(x_{k})]-f_{*}-\frac{8\theta_{3}M_{g}}{\gamma_{2}c}\right).$
Since $(1-\frac{1}{8}\gamma_{2}\alpha c)\in(0,1)$, this represents a contraction inequality. Applying the result repeatedly through iteration $k$, one obtains the desired result. ∎
When $f$ is a nonconvex function, the expected average squared norm of the gradient converges to $\left(\frac{16\gamma_{1}^{2}}{\gamma_{2}^{2}}-1\right)M_{g}$.
Theorem 3.4.
Suppose that Assumptions 3.1, 3.5 hold and that
$0<\alpha\leq\frac{\gamma_{2}}{8\gamma_{1}^{2}L},\qquad\mu_{min}\geq\frac{4\gamma_{1}\alpha}{5}.$ (35)
Then, the sequence generated by Algorithm TRishBB satisfies
$\mathbb{E}\left[\frac{1}{K}\sum_{k=1}^{K}\|\nabla f(x_{k})\|^{2}\right]\xrightarrow{K\to\infty}\left(\frac{16\gamma_{1}^{2}}{\gamma_{2}^{2}}-1\right)M_{g}.$
Proof.
Consider (30). Taking total expectation, it follows for all $k$ that
$\mathbb{E}[f(x_{k+1})]-\mathbb{E}[f(x_{k})]\leq-\frac{1}{16}\gamma_{2}\alpha\mathbb{E}[\|\nabla f(x_{k})\|^{2}]+\alpha\left(\frac{\gamma_{1}^{2}}{\gamma_{2}}-\frac{1}{16}\gamma_{2}\right)M_{g},$ (36)
which implies
$\mathbb{E}[\|\nabla f(x_{k})\|^{2}]\leq\frac{16}{\gamma_{2}\alpha}\left(\mathbb{E}[f(x_{k})]-\mathbb{E}[f(x_{k+1})]\right)+\left(\frac{16\gamma_{1}^{2}}{\gamma_{2}^{2}}-1\right)M_{g}.$ (37)
The proof is concluded by summing this inequality over all $k\in\{1,\dots,K\}$ and using the fact that $f$ is bounded below by $f_{*}$. ∎
4 Numerical Experiments
In this section, we present experimental results to compare Algorithms TRishBB_v1, TRishBB_v2 and TRishBB_v3 with the first-order version of TRish in [7], i.e., Algorithm TRishBB with $H_{k}=0$, $\forall k$. All experiments were conducted with MATLAB (R2024a) on a Rocky Linux 8.10 (64-bit) server with 256 GiB memory using an NVIDIA A100 PCIe GPU (with 80 GiB on-board memory).
4.1 Classification Problems
We consider some ML and DL test problems of the form (1). For binary classification problems using a1a, w1a, and cina datasets [36], we used logistic regression where the single loss function $f_{i}(x)$ with $x\in\mathbb{R}^{n}$ in (1) is defined as
$f_{i}(x)=\log\,(1+e^{-b_{i}(x^{T}a_{i})}),\quad i=1,\dots,N,\qquad b_{i}\in\{+1,-1\}.$
Note that $x\in\mathbb{R}^{n}$ and the data sample $a_{i}\in\mathbb{R}^{d}$ with the label $b_{i}$ have the same dimension, i.e., $n=d$. A numerically stable implementation of the logistic sigmoid was employed to prevent numerical overflow.
For multi-classification problems using MNIST [37] and CIFAR10 [38] datasets with $C$ classes, we used neural network models coupled with the softmax cross-entropy loss in (1), whose single loss function $f_{i}(x)$ is defined as
$f_{i}(x)=-\sum_{k=1}^{C}(b_{i})_{k}\log(h(a_{i};x))_{k},\quad i=1,\dots,N,$
where $a_{i}\in\mathbb{R}^{d}$ is a 3-D array (grayscale or RGB image), and the categorical label $b_{i}$ is converted into a one-hot encoded vector; i.e., it is a binary vector of length $C$, where $C-1$ elements are zero except for the element whose index corresponds to the true class that is equal to one. The output $h(a_{i};\cdot)$ is a prediction provided by the neural network (NN) whose architecture is described in Table 8 in Appendix C. The size $n$ of problem (1) using a NN depends on both the number of features $d$ and the depth of the network. We utilized the MATLAB Deep Learning Toolbox, leveraging automatic gradient computation in customized training loops; we used forward and crossentropy for implementing the NN softmax cross-entropy. See [39, Table2] for details on implementing NN classification problems by this toolbox.
More details on the test datasets and the architecture of the NNs used are available in Appendix C.
4.2 Experimental Configuration
Algorithm TRishBB_v1 was run setting $m=20$ and selecting random and independent mini-batches ${\cal{N}}_{k}$, thus the gradients $g_{k}$ are unbiased. Algorithm TRishBB_v2 was run with $\eta=0.9$ and $m=N_{b}=\lfloor N/|\mathcal{N}_{k}|\rfloor$ on the base of the discussion in Section 2.2; the mini-batches ${\cal{N}}_{k}$ were selected as disjoint subsets of ${\cal{N}}$, with ${\cal{N}}$ being shuffled at the beginning of each epoch; thus, the resulting gradients are biased. Algorithm TRishBB_v3 was run with $m_{F}=100$ as suggested in [34], its mini-batch selection was as in TRishBB_v1, and its cyclic iteration parameter $m$ was selected as in TRishBB_v2.
For all $k\geq 0$, the mini-batch size for computing stochastic gradients in these variants was $|\mathcal{N}_{k}|=64$ for the datasets a1a, w1a, and cina and $|\mathcal{N}_{k}|=128$ for the datasets MNIST and CIFAR10. The initial BB steplength $\mu_{0}$ was set to 1 in all the algorithms; moreover, the values $\mu_{\min}$, and $\mu_{\max}$ were set to $10^{-5}$ and $10^{5}$, respectively.
Regarding the predefined parameters $(\alpha,\gamma_{1},\gamma_{2})$, our settings were borrowed from those in [7] which are accompanied by a detailed discussion. The scalar $\alpha$ takes value in a prefixed set and the triplets $(\alpha,\gamma_{1},\gamma_{2})$ used in our experiments are given in Table 1 for each dataset. The parameters $\gamma_{1}$ and $\gamma_{2}$ are chosen by a gradient estimation as follows. We run the SG algorithm for one epoch with a fixed learning rate $\ell$ and compute the norm of the stochastic gradient at each iteration; then we set $G$ as the average of the norms computed throughout the epoch. The value $G$ estimates the magnitude of the stochastic gradients and promotes the occurrence of the three cases in (8) and (9), i.e., the use of normalized steps. Overall we used 60 and 36 triplets for binary classification and multi-class problems, respectively.
For the binary and multi-class test problems, given the triplet $(\alpha,\gamma_{1},\gamma_{2})$, we ran each algorithm 50 and 10 times using 50 and 10 predetermined random seeds, respectively. We present the testing accuracy, i.e., the percentage of samples in the testing set correctly identified; in case of multiple runs, the testing accuracy is averaged over the runs. In the figures shown in the following subsection, we use ExIJK to denote the results obtained in the Experiment where $\alpha$, $\gamma_{1}$, and $\gamma_{2}$ take the I-th, J-th, and K-th value indicated in Table 1, respectively.
[…]
Table 1: Hyper-parameters for the experiments.
Hyper-parameters $(G,\ell)$
ala
w1a
cina
$\alpha\in\{10^{-1},10^{-\frac{1}{2}},1,10^{\frac{1}{2}},10\}$,
$\gamma_{1}\in\left\{\frac{4}{G},\frac{8}{G},\frac{16}{G},\frac{32}{G}\right\}$, $\gamma_{2}\in\left\{\frac{1}{2G},\frac{1}{G},\frac{2}{G}\right\}$
(0.3477,0.1)(0.3477,0.1)( 0.3477 , 0.1 )
(0.0887,0.1)(0.0887,0.1)( 0.0887 , 0.1 )
(0.0497,0.1)(0.0497,0.1)( 0.0497 , 0.1 )
MNIST
CIFAR10
$\alpha\in\{10^{-3},10^{-2},10^{-1},1\}$,
$\gamma_{1}\in\left\{\frac{4}{G},\frac{8}{G},\frac{16}{G}\right\}$, $\gamma_{2}\in\left\{\frac{1}{8G},\frac{1}{4G},\frac{1}{2G}\right\}$
(0.2517,0.01)(0.2517,0.01)( 0.2517 , 0.01 )
(1.8375,0.01)(1.8375,0.01)( 1.8375 , 0.01 )
4.3 Numerical Results
We now present and analyze the numerical results obtained. Table 2 shows the percentage of unconstrained steps of the form $p_{k}=-\mu_{k}g_{k}$ taken through the iterations within five epochs; the displayed data are averaged over the runs performed. We note that in some cases this percentage is equal to zero, thus the sequence of iterates $\{x_{k}\}$ generated by our TRishBB variants coincides with the sequence generated by the first-order TRish. We also underline that the percentage of unconstrained steps increases with $\alpha$. In fact, taking into account the form of the steps (7), the larger $\alpha$ is, the higher the probability of taking unconstrained steps $(-\mu_{k}g_{k})$ is; as a consequence the performance of TRishBB variants differ considerably from the performance of TRish for large values of $\alpha$.
In Table 3, we report, for every dataset and every value of $\alpha$, the highest testing accuracy obtained by TRish and the variants of TRishBB averaging over the tested values of $\gamma_{1}$ and $\gamma_{2}$; the reported values represent the largest testing accuracies among those reached at the end of each epoch. For each algorithm we also report the ratio (”accuracy ratio”) between the smallest and the largest accuracy reported in the corresponding line of the table. We can observe that the highest testing accuracy of the TRishBB variants is comparable to or larger than the accuracy of TRish. For the datasets a1a, w1a, and cina, the values of the testing accuracy ratio are close to one for TRishBB_v2 and TRishBB_v3, namely the highest testing accuracy obtained over five epochs is not highly sensitive to the value of $\alpha$. In contrast, the testing accuracy reached by the three versions of TRishBB and TRish on the DL test problems is dependent on $\alpha$. In order to give more insight, in what follows we provide detailed statistics of the runs employing the largest value of $\alpha$ which promotes the use of the BB steplengths.
Table 2: Percentage of BB steplengths taken with different $\alpha$.
ala $\alpha=10^{-1}$ $\alpha=10^{-\frac{1}{2}}$ $\alpha=1$ $\alpha=10^{\frac{1}{2}}$ $\alpha=10\,\,$
TRishBB_v1 00.83 50.00 81.67 92.50 98.33
TRishBB_v2 03.17 67.46 88.89 96.03 99.21
TRishBB_v3 00.00 46.83 77.78 97.61 100
wla $\alpha=10^{-1}$ $\alpha=10^{-\frac{1}{2}}$ $\alpha=1$ $\alpha=10^{\frac{1}{2}}$ $\alpha=10\,\,$
TRishBB_v1 03.80 11.96 27.72 61.62 79.35
TRishBB_v2 81.96 99.48 100 100 100
TRishBB_v3 53.61 78.46 97.42 100 100
cina $\alpha=10^{-1}$ $\alpha=10^{-\frac{1}{2}}$ $\alpha=1$ $\alpha=10^{\frac{1}{2}}$ $\alpha=10\,\,$
TRishBB_v1 14.11 60.08 88.04 90.59 89.65
TRishBB_v2 71.48 99.74 100 100 100
TRishBB_v3 23.66 44.63 58.8 87.60 99.87
MNIST $\alpha=10^{-3}$ $\alpha=10^{-2}$ $\alpha=10^{-1}$ $\alpha=1$
TRishBB_v1 00.00 00.36 66.67 96.10
TRishBB_v2 00.00 00.00 00.13 87.88
TRishBB_v3 00.00 00.00 00.00 15.23
CIFAR10 $\alpha=10^{-3}$ $\alpha=10^{-2}$ $\alpha=10^{-1}$ $\alpha=1$
TRishBB_v1 00.00 00.00 24.93 81.62
TRishBB_v2 00.00 00.00 00.00 10.80
TRishBB_v3 00.00 00.00 00.00 00.78
Table 3: Highest average testing accuracy with different $\alpha$ (in %) and the accuracy ratio.
ala $\alpha=10^{-1}$ $\alpha=10^{-\frac{1}{2}}$ $\alpha=1$ $\alpha=10^{\frac{1}{2}}$ $\alpha=10$ accuracy ratio
TRishBB_v1 83.76 83.55 83.09 82.54 82.40 0.9837
TRishBB_v2 83.87 83.67 83.52 83.52 83.52 0.9958
TRishBB_v3 83.75 83.49 83.40 83.40 83.40 0.9958
TRish 83.75 83.51 81.55 79.50 79.06 0.9440
wla $\alpha=10^{-1}$ $\alpha=10^{-\frac{1}{2}}$ $\alpha=1$ $\alpha=10^{\frac{1}{2}}$ $\alpha=10$ accuracy ratio
TRishBB_v1 89.61 89.67 89.57 89.23 88.83 0.9906
TRishBB_v2 89.59 89.58 89.57 89.57 89.57 0.9998
TRishBB_v3 89.60 89.60 89.51 89.51 89.51 0.9989
TRish 89.62 89.67 89.57 89.30 88.99 0.9924
cina $\alpha=10^{-1}$ $\alpha=10^{-\frac{1}{2}}$ $\alpha=1$ $\alpha=10^{\frac{1}{2}}$ $\alpha=10$ accuracy ratio
TRishBB_v1 91.70 91.63 91.37 90.84 90.29 0.9846
TRishBB_v2 91.49 91.45 91.45 91.45 91.45 0.9996
TRishBB_v3 91.59 91.56 91.18 91.18 91.18 0.9955
TRish 91.64 91.68 90.63 87.93 87.51 0.9545
MNIST $\alpha=10^{-3}$ $\alpha=10^{-2}$ $\alpha=10^{-1}$ $\alpha=1$ accuracy ratio
TRishBB_v1 89.74 96.83 98.70 98.72 0.9090
TRishBB_v2 87.27 96.98 98.95 98.81 0.8820
TRishBB_v3 87.21 96.95 98.89 98.59 0.8819
TRish 87.21 96.95 98.87 98.42 0.8821
CIFAR10 $\alpha=10^{-3}$ $\alpha=10^{-2}$ $\alpha=10^{-1}$ $\alpha=1$ accuracy ratio
TRishBB_v1 30.86 44.72 59.19 65.41 0.4718
TRishBB_v2 31.15 44.95 60.74 65.98 0.4721
TRishBB_v3 31.09 45.06 60.28 64.89 0.4792
TRish 31.08 45.04 60.27 64.66 0.4807
4.3.1 Datasets a1a, w1a, and cina
We present the averaged accuracies obtained with $\alpha=10$ for each of the five epochs and for each triplet $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$. For this value of $\alpha$, our algorithms take the unconstrained step in the large majority of the iterations, and thus differ from TRish considerably, see Table 2.
Figure 1 illustrates the results obtained on the dataset a1a. The results displayed indicate that the highest accuracy is achieved by TRishBB_v2 and TRishBB_v3 in the first two epochs and that all TRishBB variants outperform TRish. Remarkably the testing accuracy reached by TRishBB is quite insensitive to the tested values of $\gamma_{1}$ and $\gamma_{2}$ contrary to TRish.
Figure 1: a1a: Average accuracy, $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $|\mathcal{N}_{k}|=64$. Top: TRishBB_v1, $m=20$. Middle: TRishBB_v2, $m=25$. Bottom: TRishBB_v3, $m=25$.
[…]
The results described above indicate that exploiting stochastic BB steplengths and guidelines on stochastic quasi-Newton methods improves the overall performance of TRish. Figure 4 shows the average values of the BB parameters built with $\alpha=10$ and varying $\gamma_{1}$ and $\gamma_{2}$ for each dataset. It is evident that the rule for generating $\mu_{k}$ in TRishBB_v2 and TRishBB_v3 provides smaller values of it, enhancing the selection of unconstrained steps $(-\mu_{k}g_{k})$. Algorithm TRishBB_v2 provides the best accuracy and its performance is quite insensitive to the triplet used.
Figure 2: w1a: Average accuracy, $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $|\mathcal{N}_{k}|=64$. Top: TRishBB_v1, $m=20$. Middle: TRishBB_v2, $m=38$. Bottom: TRishBB_v3, $m=38$.
Figure 3: cina: Average accuracy, $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $|\mathcal{N}_{k}|=64$. Top: TRishBB_v1, $m=20$; Middle: TRishBB_v2, $m=156$. Bottom: TRishBB_v3, $m=156$.
Figure 4: Average values of $\mu_{k}$ with triplets $(\alpha,\gamma_{1},\gamma_{2})$ along the iterations. Top: a1a, w1a and cina with $\alpha=10$. Bottom: MNIST and CIFAR10 with $\alpha=1$.
4.3.2 Datasets MNIST and CIFAR10
We present the averaged accuracies obtained for each of the five epochs performed in the experiments and for each triplet $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$. As shown in Table 2, the variants of TRishBB employed the unconstrained trust-region solution less frequently than for datasets a1a, w1a and cina, but the percentage of unconstrained steps increases with the value of $\alpha$.
Figure 5 displays the averaged accuracy obtained with the dataset MNIST. TRishBB_v2 compares well with TRishBB_v1 even though it employed a smaller percentage of unconstrained steps. Both TRishBB_v1 and TRishBB_v2 are considerably less sensitive to the values of $\gamma_{1}$ and $\gamma_{2}$ compared to TRish. TRishBB_v3 exhibits performance similar to that of TRish, which is not surprising, as it selects the unconstrained step in only approximately 15.23% of the iterations. Nevertheless, TRishBB_v3 outperforms TRish at the first epoch due to the predominance of BB steplengths in that phase; specifically, Table 4 shows the number of BB steplengths taken at the end of the first epoch compared to the total number taken at the end of the fifth epoch.
Figure 6 shows the averaged accuracies with the dataset CIFAR10. The main observation is that the proposed TRishBB variants are less sensitive than TRish to the choice of $\gamma_{1}$ and $\gamma_{2}$, see e.g., experiments Ex431, Ex432 and Ex433. Since accuracies of TRishBB variants and TRish are increasing during the epochs, we focus on the fifth epoch and observe that TRishBB_v2 performs better than all other algorithms. TRishBB_v1 and TRishBB_v3 performed better than TRish limited to some choices of $\gamma_{1}$ and $\gamma_{2}$.
Figure 4 shows that the values of $\mu_{k}$ generated by TRishBB_v1 are smaller than TRishBB_v2 and TRishBB_v3, consistently with Table 2. Although TRishBB_v1 takes many BB steplengths, TRishBB_v2 stands out as the best-performing variant as in the previous set of experiments.
Table 4: The number of BB steplengths $\mu_{k}$ taken by TRishBB_v3 on MNIST at the end of first epoch vs. the total number taken at the end of the fifth epoch.
Experiments E411 E412 E413 E421 E422 E423 E431 E432 E433
[…]
Figure 5: MNIST: Testing accuracy, $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $|\mathcal{N}_{k}|=128$. Top: TRishBB_v1, $m=20$. Middle: TRishBB_v2, $m=468$. Bottom: TRishBB_v3, $m=468$
Figure 6: CIFAR10: Testing accuracy, $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $|\mathcal{N}_{k}|=128$. Top: TRishBB_v1, $m=20$. Middle: TRishBB_v2, $m=390$. Bottom: TRishBB_v3, $m=390$.
4.3.3 Average testing loss obtained by TRishBB_v2
Sections 4.3.1 and 4.3.2 indicate that TRishBB_v2 is the overall best variant of TRishBB. In this section, we present the average testing loss vs. the number of gradient evaluations obtained with both TRishBB_v2 and TRish over five epochs. We consider specific configurations ExIJK for the triplet $(\alpha,\gamma_{1},\gamma_{2})$ described in Section 4.2. The results shown in Figure 7 refer to the highest average testing accuracy of these algorithms in Table 3 for the largest value of $\alpha$. Specifically, TRishBB_v2 achieves its best performance with configuration Ex511 on a1a, w1a, and cina, Ex422 on MNIST, and Ex413 on CIFAR10, while TRish attains its highest accuracy with Ex532 on a1a, Ex524 on w1a, Ex531 on cina, Ex411 on MNIST, and Ex421 on CIFAR10. The figure shows that our method reaches lower testing loss values earlier than TRish and that in the solution of a1a, w1a, and cina the value of the testing loss at termination is largely different and in favour of TRishBB_v2.
Figure 7: Average testing loss of TRishBB_v2 vs. number of gradient evaluations (NGE) compared to TRish across five datasets.
4.3.4 Comparison to other algorithms
As discussed in Sections 2.2 and 2.3, TRishBB_v2 borrows features from SGD-BB method given in [11], while TRishBB_v3 borrows ideas from AdaQN Algorithm given in [34]. In this section, we compare the performance of these algorithms. For TRishBB_v2 and TRishBB_v3, all hyper-parameters are set as described in Section 4.2.
TRishBB_v2 vs. SGD-BB
We implemented SGD-BB based on Algorithm 3 in [11] allowing mini-batch size greater than one. We set the cardinality of $\mathcal{N}_{k}$, the scalar $\beta$ and the cyclic update parameter $m$ as described in Section 4.2 for the experiments with TRishBB_v2.
We remark that the original SGD-BB algorithm did not apply thresholding for $\mu_{k}$ but our experiments demonstrated that thresholding is necessary for robustness, along with fine tuning for the initial value $\mu_{0}$. Specifically, initially, we ran SGD-BB with $\mu_{0}=1$ which was the setting used in TRishBB_v2 but the BB steplength generated in SGD-BB were too large causing failures in the fist epoch. To mitigate this issue, we ran the algorithm with $\mu_{0}=10^{-2}$ for a1a, w1a, and cina, and with $\mu_{0}=10^{-3}$ for MNIST and CIFAR10. Nonetheless, these smaller values for $\mu_{0}$ were still insufficient to provide good results in the next epochs. Hence, we applied the thresholding of $\mu_{k}$ as in TRishBB_v2 and verified that setting $\mu_{\min}=10^{-5}$ and $\mu_{\max}=10^{-1}$ provided convergence for all datasets.
Table 5 compares the performance of SGD-BB against TRishBB_v2 performed with the largest values of $\alpha$ indicated in Sections 4.3.1 and 4.3.2. The highest average testing accuracy achieved by TRishBB_v2 per epoch is close to or better than the accuracy achieved by tuned SGD-BB. Indeed, in SGD-BB a careful adjustment of both the initial stepsize $\mu_{0}$ and the threshold $\mu_{\max}$ was necessary to obtain good performance and the average testing accuracy may differ greatly from one epoch to the other.
Table 5: Highest average testing accuracy per epoch obtained by TRishBB_v2 with the largest $\alpha$ and by SGD-BB.
[…]
TRishBB_v2 83.14 83.52 79.71 80.94 80.77
SGD-BB 82.63 81.73 60.03 61.76 54.57
[…]
TRishBB_v2 89.38 89.44 89.47 89.53 89.57
SGD-BB 89.38 89.46 89.54 89.45 87.74
[…]
TRishBB_v2 90.07 90.99 91.35 91.45 91.20
SGD-BB 89.52 91.01 91.32 83.99 49.21
[…]
TRishBB_v2 98.30 98.61 98.73 98.71 98.81
SGD-BB 93.44 95.01 73.08 98.26 98.70
[…]
TRishBB_v2 48.18 58.42 62.52 65.01 65.98
SGD-BB 20.54 27.49 19.57 51.67 57.83
TRishBB_v3 vs. AdaQN
Our experiments with AdaQN were performed using adaQN.m from the package minSQN, available at https://github.com/keskarnitish/minSQN/blob/master/%2Bmethods/adaQN.m. As suggested therein and in the reference paper [34], we set the hyper-parameters
We set the limited-memory size of the L-BFGS Hessian approximations $H_{k}$ to 202020, and the Fisher information storage size $m_{F}$ to 100100100. AdaQN provides two strategies for initializing $H_{0}$. We adopted the default option, RMSprop-like initialization, which showed better performance in our experiments compared to the Adagrad-like strategy.
The cyclic iteration parameter $L$ in AdaQN, corresponding to our parameter $m$ in TRishBB_v3, needs fine tuning and we observed that it is highly sensitive to the choice of random seed during hyper-parameter tuning. We fixed its value in the set $L\in\{2,5,10,20,64,m\}$ where the parameter $m$ is the same as in TRishBB_v3. The remaining hyper-parameter values were adopted from [34] and the implementation mentioned above. Following this implementation, the stepsize $\alpha$ in AdaQN was sampled from a log-uniform distribution over the interval $[10^{-6},10^{2}]$. We tuned AdaQN using 20 different random seeds for the values of $L$ and $\alpha$, as part of the standard procedure in the adaQN.m routine, with each run lasting 5 epochs. Using the resulting optimal hyper-parameters $L$ and $\alpha$, we obtained the results presented in Table 6 where AdaQN is compared with TRishBB_v3 using the largest $\alpha$ indicated in Sections 4.3.1 and 4.3.2.
Table 6: Highest average testing accuracy per epoch obtained by TRishBB_v3 with the largest $\alpha$ and by AdaQN with its tuned $(L,\alpha)$.
[…]
TRishBB_v3 83.19 83.40 82.60 78.65 79.42
AdaQN $(L=10,\,\alpha=1.13\times 10^{-3})$ 76.62 78.83 80.34 80.52 80.48
[…]
TRishBB_v3 89.34 89.42 89.51 89.47 89.24
AdaQN ($L=20,\,\alpha=8.41\times 10^{-3}$) 48.27 47.78 47.65 47.70 47.57
[…]
TRishBB_v3 90.00 91.18 90.07 88.58 85.05
AdaQN ($L=5,\,\alpha=7.21\times 10^{-3}$) 88.73 87.76 87.92 88.25 88.14
[…]
TRishBB_v3 98.28 97.91 98.26 98.25 98.59
AdaQN ($L=468,\,\alpha=6.50\times 10^{-4}$) 98.13 98.34 98.66 98.82 98.86
[…]
TRishBB_v3 49.55 58.45 61.78 63.80 64.89
AdaQN ($L=64,\,\alpha=4.73\times 10^{-4}$) 54.65 60.65 62.99 66.02 66.46
The table shows that the highest average testing accuracy achieved by TRishBB_v3 per epoch over a1a, w1a, and cina is generally larger than the accuracy achieved by AdaQN, in particular for earlier epochs. On the other hand, the performance of both TRishBB_v3 and AdaQN is close to each other for MNIST while AdaQN outperforms TRishBB_v3 in the solution of CIFAR10. In MNIST and CIFAR10 problems, TRishBB_v3 could meet rarely the iterations at which the BB steplength is computed, see Table 2; in addition, the good performance of AdaQN is motivated by the use of an L-BFGS-based Hessian approximation coupled with a step acceptance criterion based on sampled function evaluations.
4.3.5 The cyclic iteration parameter $m$ in TRishBB_v3
We conclude our experimental analysis, focusing on the cyclic iteration parameter $m$ in TRishBB_v3. In all experiments conducted we have fixed $m=N_{b}$ with $N_{b}=\lfloor N/|\mathcal{N}_{k}|\rfloor$ ($N_{b}$ equal to 25, 38, 156, 468, and 390 for a1a, w1a, cina, MNIST, and CIFAR10, respectively). To assess the sensitivity of TRishBB_v3 to the hyper-parameter $m$, we report in Figures 89 additional results obtained with $m\in\{5,20,64,N_{b}\}$, i.e., the values used for tuning Ada-QN; to improve readability in the figures we have eliminated the results corresponding to $m=2$ and $m=10$ because $m=5$ sufficiently represents the impact of small values of $m$. These figures show that for large values of $m$, $m=64$ and $m=N_{b}$, the performance is not highly sensitive to the choice of $\gamma_{1}$ and $\gamma_{2}$ and that the setting $m=N_{b}$ provides high accuracy at some epochs for all tests, making unnecessary to tune a further hyper-parameter.
5 Conclusion
We introduced the Algorithm TRishBB applicable to finite-sum minimization problems. It is a stochastic trust-region method that combines the TRish methodology with stochastic Barzilai-Borwein steplengths to inject cheap second-order information. We presented theoretical analysis of TRishBB and devised three practical variants differing in the computation of the BB steplength. TRishBB_v1 mimics the implementation of a standard BB rule, TRishBB_v2 employs an accumulated BB parameter per iteration, TRishBB_v3 retrieves the second-order information by the accumulated Fisher Information matrix. Numerical results demonstrated that, for large values of $\alpha$, TRishBB_v2 and TRishBB_v3 are quite insensitive to the hyper-parameters and performs better than the first-order TRish method. We also observe that they compare well with two stochastic quasi-Newton methods from the literature. Taking into account that TRishBB_v3 is applicable as long as it is feasible the use of the empirical Fischer Information, TRishBB_v2 appears to be a robust and general algorithm.
Regarding future investigations, it is worthy to further explore the definition of stochastic BB parameters, including stochastic diagonal BB rules, as suggested by [41].
Figure 8: Testing average accuracy of TRishBB_v3 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m\in\{5,20,64,N_{b}\}$. Top: a1a, $N_{b}=25$. Middle: w1a, $N_{b}=38$. Bottom: cina, $N_{b}=156$.
Figure 9: Testing average accuracy of TRishBB_v3 with $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $m\in\{5,20,64,N_{b}\}$. Top: MNSIT, $N_{b}=468$. Bottom: CIFAR10, $N_{b}=390$.
A Some intermediate results
We present some theoretical results required in Appendix B.
Lemma A.1.
Suppose Assumption 3.1 holds, then
$f(x_{k+1})-f(x_{k})\leq\underbrace{g_{k}^{T}p_{k}+\frac{1}{2\mu_{k}}\|p_{k}\|^{2}}_{m_{k}(p_{k})}+(\nabla f(x_{k})-g_{k})^{T}p_{k}+\frac{L}{2}\|p_{k}\|^{2}.$ (38)
Proof.
By (10), we have
$f(x_{k+1})-f(x_{k})\leq\nabla f(x_{k})^{T}p_{k}+\frac{L}{2}\|p_{k}\|^{2}.$
Adding, subtracting $g_{k}^{T}p_{k}$, and using the fact that $\frac{1}{2\mu_{k}}\|p_{k}\|^{2}\geq 0$ conclude the proof. ∎
Lemma A.2.
Suppose Assumption 3.1 holds, then
$m_{k}(p_{k})=g_{k}^{T}p_{k}+\frac{1}{2\mu_{k}}\|p_{k}\|^{2}\leq-\Delta_{k}\|g_{k}\|+\frac{1}{2\mu_{k}}\Delta_{k}^{2}.$ (39)
Proof.
If $p_{k}=-\mu_{k}g_{k}$, then $m_{k}(p_{k})=-\frac{1}{2}\mu_{k}\|g_{k}\|^{2}$. If $p_{k}=-\frac{\Delta_{k}}{\|g_{k}\|}g_{k}$, then $m_{k}(p_{k})=-\Delta_{k}\|g_{k}\|+\frac{1}{2\mu_{k}}\Delta_{k}^{2}$. Since $m_{k}(-\mu g_{k})$ is the minimum value attainable by $m_{k}$, the proof is concluded. ∎
B Proofs from Section 3
B.1 Proof of Lemma 3.1
Proof.
Given the step taken by the Algorithm 1, i.e., $p_{k}=x_{k+1}-x_{k}$, the inequality (10) implies
$f(x_{k+1})=f(x_{k}+p_{k})\leq f(x_{k})+\nabla f(x_{k})^{T}p_{k}+\frac{L}{2}\|p_{k}\|^{2}.$ (40)
Thus, considering the events that Case 1–3 occur, i.e., $C_{i,k}$ for $i=1,2,3$, the law of total probability gives
$\mathbb{E}_{k}[f(x_{k+1})]-f(x_{k})\leq\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}]+\frac{L}{2}\mathbb{E}_{k}[\|p_{k}\|^{2}]$
$=$ $\sum_{i=1}^{3}\mathbb{P}_{k}[C_{i,k}]\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{i,k}]+\frac{L}{2}\sum_{i=1}^{3}\mathbb{P}_{k}[C_{i,k}]\mathbb{E}_{k}[\|p_{k}\|^{2}|C_{i,k}].$ (41)
Further,
$\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{i,k}]$ $=$ $\mathbb{P}_{k}[\overline{S_{BB}}|C_{i,k}]\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{i,k}\cap\overline{S_{BB}}]$ (42)
$+$ $\mathbb{P}_{k}[S_{BB}|C_{i,k}]\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{i,k}\cap S_{BB}],$
and
$\mathbb{E}_{k}[\|p_{k}\|^{2}|C_{i,k}]$ $=$ $\mathbb{P}_{k}[\overline{S_{BB}}|C_{i,k}]\mathbb{E}_{k}[\|p_{k}\|^{2}|C_{i,k}\cap\overline{S_{BB}}]$ (43)
$+$ $\mathbb{P}_{k}[S_{BB}|C_{i,k}]\mathbb{E}_{k}[\|p_{k}\|^{2}|C_{i,k}\cap S_{BB}],$
for $i=1,2,3$. Now, we analyze the events separately.
Event ${C_{1,k}}$. Consider the event $\overline{C_{1,k}^{BB}}\stackrel{{\scriptstyle\rm def}}{{=}}C_{1,k}\cap\overline{S_{BB}}$. It holds $p_{k}=-\gamma_{1}\alpha g_{k}$, and [7, Equations (8), (9)] gives
$\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|\overline{C_{1,k}^{BB}}]$ $\leq\ -\gamma_{1}\alpha\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}g_{k}|\overline{C_{1,k}^{BB}}\right]$
$+(\gamma_{1}-\gamma_{2})\alpha\mathbb{P}_{k}\left[{\cal{E}}_{k}|\overline{C_{1,k}^{BB}}\right]\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}g_{k}|\overline{C_{1,k}^{BB}}\cap{\cal{E}}_{k}\right],$
and
$\mathbb{E}_{k}\left[\|p_{k}\|^{2}|\overline{C_{1,k}^{BB}}\right]=\gamma_{1}^{2}\alpha^{2}\mathbb{E}_{k}\left[\|g_{k}\|^{2}|\overline{C_{1,k}^{BB}}\right].$
In the event $C_{1,k}^{BB}\stackrel{{\scriptstyle\rm def}}{{=}}C_{1,k}\cap S_{BB}$, it holds $p_{k}=-\mu_{k}g_{k}$ with $\mu_{k}\leq\gamma_{1}\alpha$, and paralleling [7, Equations (8) and (9)] results in
$\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{1,k}^{BB}]$ $=\mathbb{E}_{k}[-\mu_{k}\nabla f(x_{k})^{T}g_{k}|C_{1,k}^{BB}]$
$\leq(\gamma_{1}\alpha-\mu_{min})\mathbb{P}_{k}[{\cal{E}}_{k}|C_{1,k}^{BB}]\mathbb{E}_{k}[\nabla f(x_{k})^{T}g_{k}|C_{1,k}^{BB}\cap{\cal{E}}_{k}]$
$-\gamma_{1}\alpha\mathbb{E}_{k}[\nabla f(x_{k})^{T}g_{k}|C_{1,k}^{BB}],$
and
$\mathbb{E}_{k}[\|p_{k}\|^{2}|C_{1,k}^{BB}]\leq\gamma_{1}^{2}\alpha^{2}\mathbb{E}_{k}[\|g_{k}\|^{2}|C_{1,k}^{BB}].$
Now, we combine (42), (43) and the previous results. We take into account that $(\gamma_{1}-\gamma_{2})\alpha\leq(\gamma_{1}\alpha-\mu_{min})$ by Assumption 3.4 and
$\mathbb{P}_{k}({\cal{E}}_{k}|\overline{C_{1,k}^{BB}})\leq\mathbb{P}_{k}({\cal{E}}_{k}|C_{1,k}),\qquad\mathbb{P}_{k}({\cal{E}}_{k}|C_{1,k}^{BB})\leq\mathbb{P}_{k}({\cal{E}}_{k}|C_{1,k}).$
Thus, we get
$\,\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{1,k}]\leq\,(\gamma_{1}\alpha-\mu_{min})\mathbb{P}_{k}[S_{BB}|C_{1,k}]\mathbb{P}_{k}({\cal{E}}_{k}|C_{1,k}^{BB})\mathbb{E}_{k}[\nabla f(x_{k})^{T}g_{k}|C_{1,k}^{BB}\cap{\cal{E}}_{k}]\Bigr{)}$
[…]
i.e.,
$\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{1,k}]$ $\leq(\gamma_{1}\alpha-\mu_{min})\mathbb{P}_{k}({\cal{E}}_{k}|C_{1,k})\mathbb{E}_{k}[\nabla f(x_{k})^{T}g_{k}|C_{1,k}\cap{\cal{E}}_{k}]$
$-\gamma_{1}\alpha\mathbb{E}_{k}[\nabla f(x_{k})^{T}g_{k}|C_{1,k}],$
and
$\mathbb{E}_{k}[\|p_{k}\|^{2}|C_{1,k}]\leq\gamma_{1}^{2}\alpha^{2}\mathbb{E}_{k}[\|g_{k}\|^{2}|C_{1,k}].$ (44)
Event ${C_{2,k}}$. The event $C_{2,k}$ occurs when $\|g_{k}\|^{-1}\leq\gamma_{1}$ and $\|g_{k}\|^{-1}\geq\gamma_{2}$. In the event $\overline{C_{2,k}^{BB}}\stackrel{{\scriptstyle\rm def}}{{=}}C_{2,k}\cap\overline{S_{BB}}$, it holds $p_{k}=-\alpha g_{k}/\|g_{k}\|$ and [7, Equations (10) and (11)] gives
$\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|\overline{C_{2,k}^{BB}}]$ $\leq-\gamma_{1}\alpha\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}g_{k}|\overline{C_{2,k}^{BB}}\right]$
$+(\gamma_{1}-\gamma_{2})\alpha\mathbb{P}_{k}\left[{\cal{E}}_{k}|\overline{C_{2,k}^{BB}}\right]\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}g_{k}|\overline{C_{2,k}^{BB}}\cap{\cal{E}}_{k}\right]$
and
$\mathbb{E}_{k}\left[\|p_{k}\|^{2}|\overline{C_{2,k}^{BB}}\right]=\alpha^{2}\leq\gamma_{1}^{2}\alpha^{2}\mathbb{E}_{k}\left[\|g_{k}\|^{2}|\overline{C_{2,k}^{BB}}\right].$
Assumption 3.4 ensures that $\mu_{min}\leq\frac{\alpha}{\|g_{k}\|}$ and in the event $C_{2,k}^{BB}\stackrel{{\scriptstyle\rm def}}{{=}}C_{2,k}\cap S_{BB}$, it holds $\mu_{min}\leq\mu_{k}\leq\frac{\alpha}{\|g_{k}\|}$. Then,
$\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}p_{k}|C_{2,k}^{BB}\right]$ $\leq-\mu_{min}\mathbb{P}_{k}\left[{\cal{E}}_{k}|C_{2,k}^{BB}\right]\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}g_{k}\bigg{|}C_{2,k}^{BB}\cap{\cal{E}}_{k}\right]$
$-\alpha\mathbb{P}_{k}\left[\overline{{\cal{E}}_{k}}|C_{2,k}^{BB}\right]\mathbb{E}_{k}\left[\frac{\nabla f(x_{k})^{T}g_{k}}{\|g_{k}\|}\bigg{|}C_{2,k}^{BB}\cap\overline{{\cal{E}}_{k}}\right]$
$\leq-\mu_{min}\mathbb{P}_{k}\left[{\cal{E}}_{k}|C_{2,k}^{BB}\right]\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}g_{k}|C_{2,k}^{BB}\cap{\cal{E}}_{k}\right]$
$-\gamma_{1}\alpha\mathbb{P}_{k}\left[\overline{{\cal{E}}_{k}}|C_{2,k}^{BB}\right]\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}g_{k}|C_{2,k}^{BB}\cap\overline{{\cal{E}}_{k}}\right]$
$=-\gamma_{1}\alpha\mathbb{E}_{k}[\nabla f(x_{k})^{T}g_{k}|C_{2,k}^{BB}]$
$+(\gamma_{1}\alpha-\mu_{min})\mathbb{P}_{k}\left[{\cal{E}}_{k}|C_{2,k}^{BB}\right]\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}g_{k}|C_{2,k}^{BB}\cap{\cal{E}}_{k}\right],$
and
$\mathbb{E}_{k}[\|p_{k}\|^{2}|C_{2,k}^{BB}]\leq\alpha^{2}\leq\gamma_{1}^{2}\alpha^{2}\mathbb{E}_{k}[\|g_{k}\|^{2}|C_{2,k}^{BB}].$
Now, we combine (42), (43) and the previous results. We take into account that $(\gamma_{1}-\gamma_{2})\alpha\leq(\gamma_{1}\alpha-\mu_{min})$ by Assumption 3.4 and that
$\mathbb{P}_{k}\left({\cal{E}}_{k}|\overline{C_{2,k}^{BB}}\right)\leq\mathbb{P}_{k}\left({\cal{E}}_{k}|C_{2,k}\right),\qquad\mathbb{P}_{k}\left({\cal{E}}_{k}|C_{2,k}^{BB}\right)\leq\mathbb{P}_{k}\left({\cal{E}}_{k}|C_{2,k}\right).$
It follows
$\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{2,k}]\leq-\gamma_{1}\alpha\Bigl{(}\mathbb{P}_{k}[\overline{S_{BB}}|C_{2,k}]\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}p_{k}|\overline{C_{2,k}^{BB}}\right]$
$+\mathbb{P}_{k}\left[S_{BB}|C_{2,k}\right]\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}p_{k}|C_{2,k}^{BB}\right]\Bigr{)}$
$+(\gamma_{1}\alpha-\mu_{min})\Bigl{(}\mathbb{P}_{k}\left[\overline{S_{BB}}|C_{2,k}\right]\mathbb{P}_{k}\left[{\cal{E}}_{k}|C_{2,k}\cap\overline{S_{BB}}\right]\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}p_{k}|\overline{C_{2,k}^{BB}}\cap{\cal{E}}_{k}\right]$
$+\mathbb{P}_{k}[S_{BB}|C_{2,k}]\mathbb{P}_{k}\left[{\cal{E}}_{k}|C_{2,k}^{BB}\right]\mathbb{E}_{k}\left[\nabla f(x_{k})^{T}p_{k}|C_{2,k}^{BB}\cap{\cal{E}}_{k}\right]\Bigr{)},$
and finally
$\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{2,k}]\leq$ $(\gamma_{1}\alpha-\mu_{min})\mathbb{P}_{k}({\cal{E}}_{k}|C_{2,k})\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{2,k}\cap{\cal{E}}_{k}]$
-- $\gamma_{1}\alpha\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{2,k}].$ (45)
Further,
$\mathbb{E}_{k}[\|p_{k}\|^{2}|C_{2,k}]\leq\alpha^{2}\leq\gamma_{1}^{2}\alpha^{2}\mathbb{E}_{k}[\|g_{k}\|^{2}|C_{2,k}].$ (46)
Event ${C_{3,k}}$. Following [7, Equations (12) and(13)] and the lines of the proof for the event $C_{1,k}$, we obtain
$\mathbb{E}_{k}[\nabla f(x_{k})^{T}p_{k}|C_{3,k}]\leq$ $(\gamma_{1}\alpha-\mu_{min})\mathbb{P}_{k}({\cal{E}}_{k}|C_{3,k})\mathbb{E}_{k}[\nabla f(x_{k})^{T}g_{k}|C_{3,k}\cap{\cal{E}}_{k}]$
-- $\gamma_{1}\alpha\mathbb{E}_{k}[\nabla f(x_{k})^{T}g_{k}|C_{3,k}],$ (47)
and
$\mathbb{E}_{k}[\|p_{k}\|^{2}|C_{3,k}]\leq\gamma_{1}^{2}\alpha^{2}\mathbb{E}_{k}[\|g_{k}\|^{2}|C_{3,k}].$ (48)
Combining (41), (47) and (48), the claim follows. ∎
B.2 Proof of Lemma 3.3
Proof.
Regarding (29), we analyze the different events ${C_{1,k}}$, $i=1,2,3$, separately.
Event ${C_{1,k}}$. Since $\Delta_{k}=\gamma_{1}\alpha\|g_{k}\|$, the rightmost inequality in (35) and (39) give
$m_{k}(p_{k})\leq-\gamma_{1}\alpha\left(1-\frac{1}{2}\frac{\gamma_{1}\alpha}{\mu_{k}}\right)\|g_{k}\|^{2}\leq-\frac{3}{8}\gamma_{1}\alpha\|g_{k}\|^{2}.$
Using $\|p_{k}\|\leq\gamma_{1}\alpha\|g_{k}\|$, Lemma A.1, the leftmost inequality in (35) and $\gamma_{1}\geq\gamma_{2}$, we find that
$\ f(x_{k+1})-f(x_{k})$
$\leq-\frac{3}{8}\gamma_{1}\alpha\|g_{k}\|^{2}+\gamma_{1}\alpha\|\nabla f(x_{k})-g_{k}\|\|g_{k}\|+\frac{1}{2}L\gamma_{1}^{2}\alpha^{2}\|g_{k}\|^{2}$
$\leq-\frac{3}{8}\gamma_{1}\alpha\|g_{k}\|^{2}+\gamma_{1}\alpha\|\nabla f(x_{k})-g_{k}\|\|g_{k}\|+\frac{1}{16}\gamma_{1}\alpha\|g_{k}\|^{2}.$
Then, by $\gamma_{1}\geq\gamma_{2}$ and
$0\leq\left(\frac{1}{2}\|g_{k}\|-\|\nabla f(x_{k})-g_{k}\|\right)^{2}=\tfrac{1}{4}\|g_{k}\|^{2}-\|\nabla f(x_{k})-g_{k}\|\|g_{k}\|+\|\nabla f(x_{k})-g_{k}\|^{2},$
we obtain
$\ f(x_{k+1})-f(x_{k})$ (49)
$\leq-\frac{1}{16}\gamma_{1}\alpha\|g_{k}\|^{2}+\gamma_{1}\alpha\|\nabla f(x_{k})-g_{k}\|^{2}$
$\leq-\frac{1}{16}\gamma_{2}\alpha\|g_{k}\|^{2}+\frac{\gamma_{1}^{2}}{\gamma_{2}}\alpha\|\nabla f(x_{k})-g_{k}\|^{2}.$
Event ${C_{2,k}}$. Since $\Delta_{k}=\alpha$, using (39), the rightmost inequality in (35) and $1/\gamma_{1}\leq\|g_{k}\|\leq 1/\gamma_{2}$ gives
$m_{k}(p_{k})\leq-\alpha\|g_{k}\|+\frac{1}{2}\frac{\alpha^{2}}{\mu_{k}}\leq-\alpha\|g_{k}\|+\frac{5}{8}\frac{\alpha}{\gamma_{1}}\leq-\frac{3}{8}\alpha\|g_{k}\|\leq-\frac{3}{8}\gamma_{2}\alpha\|g_{k}\|^{2}.$
Using $1/\gamma_{1}\leq\|g_{k}\|\leq 1/\gamma_{2}$, Lemma A.1, and the leftmost inequality in (35), we find that
$\ f(x_{k+1})-f(x_{k})\leq$ $\ -\frac{3}{8}\gamma_{2}\alpha\|g_{k}\|^{2}+\|\nabla f(x_{k})-g_{k}\|\alpha+\frac{1}{2}L\alpha^{2}$
$\leq$ $\ -\frac{3}{8}\gamma_{2}\alpha\|g_{k}\|^{2}+\ \|\nabla f(x_{k})-g_{k}\|\alpha+\frac{1}{2}\gamma_{1}^{2}\alpha^{2}L\|g_{k}\|^{2}$
$\leq$ $\ -\frac{3}{8}\gamma_{2}\alpha\|g_{k}\|^{2}+\ \|\nabla f(x_{k})-g_{k}\|\alpha+\frac{1}{16}\gamma_{2}\alpha\|g_{k}\|^{2}.$
Since
$0\leq\frac{\gamma_{2}}{\gamma_{1}^{2}}\left(\frac{1}{2}-\frac{\gamma_{1}^{2}}{\gamma_{2}}\|\nabla f(x_{k})-g_{k}\|\right)^{2}=\frac{\gamma_{2}}{4\gamma_{1}^{2}}-\|\nabla f(x_{k})-g_{k}\|+\frac{\gamma_{1}^{2}}{\gamma_{2}}\|\nabla f(x_{k})-g_{k}\|^{2},$
we obtain
$f(x_{k+1})-f(x_{k})\leq\ -\frac{1}{16}\gamma_{2}\alpha\|g_{k}\|^{2}+\frac{\gamma_{1}^{2}}{\gamma_{2}}\alpha\|\nabla f(x_{k})-g_{k}\|^{2}.$
Event ${C_{3,k}}$. The proof trivially follows what we argued for the event ${C_{1,k}}$ as we have $\gamma_{1}\geq\gamma_{2}$.
Summarizing the analysis above, inequality (29) holds. Finally, regarding (30), due to (26) and (27), inequality (29) yields
$\ \mathbb{E}_{k}[f(x_{k+1})]-f(x_{k})$
$\leq$ $\ -\frac{1}{16}\gamma_{2}\alpha(\|\nabla f(x_{k})\|^{2}+\mathbb{E}_{k}[\|\nabla f(x_{k})-g_{k}\|^{2}])+\frac{\gamma_{1}^{2}}{\gamma_{2}}\alpha\mathbb{E}_{k}[\|\nabla f(x_{k})-g_{k}\|^{2}]$
$=$ $\ -\frac{1}{16}\gamma_{2}\alpha\|\nabla f(x_{k})\|^{2}+\alpha\left(\frac{\gamma_{1}^{2}}{\gamma_{2}}-\frac{1}{16}\gamma_{2}\right)\mathbb{E}_{k}[\|\nabla f(x_{k})-g_{k}\|^{2}]$
which give the thesis. ∎
C Classification Problems
Table 7: Classification test problems.
Dataset $\mathbf{(N,{N}_{t})}$ $\mathbf{d}$ $\mathbf{n}$ Type $\mathbf{f_{i}(x)}$ NN
a1a [36] (1605,29351)(1605,29351)( 1605 , 29351 ) 123123123 123123123 Binary Logistic regression -
w1a [36] (2477,47272)(2477,47272)( 2477 , 47272 ) 300300300 300300300 Binary Logistic regression -
cina [36] (10000,6033)(10000,6033)( 10000 , 6033 ) 132132132 132132132 Binary Logistic regression -
MNIST [37] (60000,10000)(60000,10000)( 60000 , 10000 ) 28×28×128$\times$ 28$\times$ 128 × 28 × 1 431030431030431030 Multi Cross-entropy NET-1
CIFAR10 [38] (50000,10000)(50000,10000)( 50000 , 10000 ) 32×32×332$\times$ 32$\times$ 332 × 32 × 3 545098545098545098 Multi Cross-entropy NET-2
Table 7 summarizes the five datasets used in our experiments for both binary and multi-class classification tasks in Section 4.1. Each test dataset with $C$ possible classes is divided into training and testing sets, denoted as $\{(a_{i},b_{i})\,|i\in\mathcal{N}\}$ and $\{(a_{i},b_{i})\,|i\in\mathcal{N}_{t}\}$, with $|\mathcal{N}|=N$ and $|\mathcal{N}_{t}|={N}_{t}$ samples, respectively; the values of $N$ and ${N}_{t}$ are reported in the table. The vector $a_{i}\in\mathbb{R}^{d}$ represents the $i$-th data with $d$ features while $b_{i}$ is the corresponding categorical label associated.
The dataset cina was normalized such that each feature vector has components in the interval [0,1][0,1][ 0 , 1 ]. For MNIST and CIFAR10, besides the zero-one rescaling, we implemented z-score normalization to have zero mean and unit variance. Precisely, all $N$ training images, (3-D arrays of size $\sigma_{1}\times\sigma_{2}\times\sigma_{2}$ where $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$ denote the height, width, and number of channels of the images, respectively), undergo normalization by subtracting the mean (a 3-D array) and dividing by the standard deviation (a 3-D array) component-wise. Data in the testing set are also normalized using the same statistics as in the training data. The architecture of the neural network (NN) is described in Table 8.111The compound $Conv(5,20,1,0){/}ReLU{/}MaxPool(2,2,0)$ indicates a convolutional layer ($Conv$) using 202020 filters of size $\times 5$ 5, stride 111, and no padding, a nonlinear activation function ($ReLu$) and, a 2-D max-pooling layer with a channel of size $\times 2$ 2, stride 2 and no padding. The syntax $FC(C){/}Softmax$ denotes a fully connected layer of $C$ neurons followed by the $softmax$ layer. For MNIST dataset with grayscaled images in $C=10$ classes (digits 0–9), and CIFAR10 with RGB images in $C=10$ categories (e.g., dogs and ships), we used NET-1 and NET-2 respectively.
Table 8: NN architecture.
NET-1 $Input(d)$
$Conv(5,20,1,0){/}ReLU{/}MaxPool(2\times 2,2,0)$
$Conv(5,50,1,0){/}ReLU{/}MaxPool(2\times 2,2,0)$
$FC(500){/}ReLU$
$FC(C){/}Softmax$
NET-2 $Input(d)$
$Conv(3,32,1,same){/}ReLU{/}AveragePool(2\times 2,2,0)$
$Conv(3,64,1,same){/}ReLU{/}AveragePool(2\times 2,2,0)$
$FC(128){/}ReLU$
$FC(C){/}Softmax$
Acknowledgments
The authors are members of the INdAM-GNCS Research Group. This research was partially supported by INDAM-GNCS through Progetti di Ricerca 2023, by PNRR - Missione 4 Istruzione e Ricerca - Componente C2 Investimento 1.1, Fondo per il Programma Nazionale di Ricerca e Progetti di Rilevante Interesse Nazionale (PRIN) funded by the European Commission under the NextGeneration EU programme, project “Advanced optimization METhods for automated central veIn Sign detection in multiple sclerosis from magneTic resonAnce imaging (AMETISTA)”, code: P2022J9SNP, MUR D.D. financing decree n. 1379 of 1st September 2023 (CUP E53D23017980001), project “Numerical Optimization with Adaptive Accuracy and Applications to Machine Learning”, code: 2022N3ZNAXMUR D.D. financing decree n. 973 of 30th June 2023 (CUP B53D23012670006), and by Partenariato esteso FAIR “Future Artificial Intelligence Research” SPOKE 1 Human-Centered AI. Obiettivo 4, Project “Mathematical and Physical approaches to innovative Machine Learning technologies (MaPLe)”.
Competing interests declarations
The authors have no relevant interests to disclose.
Data availability statements
The datasets utilized in this research are publicly accessible and commonly employed benchmarks in the field of Machine Learning and Deep Learning, see [36, 37].